trains A and B are traveling in the same direction on parallel tracks.train A is traveling at 40 miles per hour and train B is traveling at 60 miles per hour . train A passes a station at 9:20 am if train B passes the same station at 9:35 Am at what time will train B catch up to train A. can you please help me salve this quition?
Found 2 solutions by lwsshak3, mananth:
Answer by lwsshak3(11628)

trains A and B are traveling in the same direction on parallel tracks.train A is traveling at 40 miles per hour and train B is traveling at 60 miles per hour . train A passes a station at 9:20 am if train B passes the same station at 9:35 Am at what time will train B catch up to train A. can you please help me salve this quition?
..
Let x=hours train B will catch up with train A
When train B catches up, train A travel time will be 15 minutes or 1/4 hour longer than travel time of train B.
That is, x+1/4 = hours train A traveled
Distance=speed*travel time
60x=40(x+1/4)
60x=40x+10
20x=10
x=1/2 hour
ans:
Train B will catch up with train A at 9:50 AM, 30 minutes after passing the station

Answer by mananth(16946)

trains A and B are traveling in the same direction on parallel tracks.train A is traveling at 40 miles per hour and train B is traveling at 60 miles per hour . train A passes a station at 9:20 am if train B passes the same station at 9:35 Am at what time will train B catch up to train A. can you please help me salve this quition?
Train A speed 40mph
Train B speed 60mph

Train A passes station at 09:20
Train B passes station at 09:35

Difference in time=00:15 =15minutes
Train A will have covered 10miles before Train B starts

catch up distance=10miles
catch up speed = 60-40 mph
catch up speed = 20 mph

Catchup time = catchup distance/catch up speed
catch up time= 10/20
catch up time= 0.5hours
catch up time= 30 minutes
Train will catch up at 10:05 AM